Fermi National Laboratory
 

Accelerator Update for February 17, 2006 - February 20, 2006

 

Friday February 17

The day shift began with Operations monitoring the special 1x1 low luminosity store and the Recycler stash, with the Collider Detector at Fermilab (CDF) and D-Zero (D0) using the special store, with the Antiproton Source (Pbar) conducting reverse proton studies, and with the Booster and Linac in access.

The Booster technicians completed their work and came out of access at 8:30 AM.

Tevatron (TeV) experts began an end-of-store study at 8:40 AM.

The Linac work ended, Operations established beam, and an expert tuned the beam line at 9:36 AM.

At 9:44 AM, Operations reported that two Linac RF stations (LRF1 & 2) had been tripping off every few minutes. Operators had been able to reset the stations after each trip.

Operations terminated the special 1x1 low luminosity store at 9:59 AM.

Operations established beam to MiniBooNE at 10:02 AM.

Operations established beam to NuMI at 10:11 AM.

The end-of-store study ended and Operations prepared D0 for an access at 10:14 AM.

Pbar studies ended and Pbar resumed stacking at 10:20 AM.

Operations put the TeV through a dry squeeze around 11:50 AM.

Operations put the TeV into shot setup at 1:05 PM.

At 2:48 PM, Operations reported that the TeV suffered a sector A1 quench during the squeeze. TeV experts investigated.

A Main Injector (MI) expert began a slip stacking study at 3:21 PM.

Operations began preparing the Recycler to stash antiprotons at 7:18 PM.

Operations put the TeV through a dry squeeze at 7:26 PM.

The stash ended and Pbar resumed stacking at 8:06 PM.

Saturday February 18

The midnight shift began with Operations monitoring the Recycler stash, with Pbar stacking, and with NuMI, MiniBooNE, MTest, and MCenter taking beam.

Operations put the TeV through a wet squeeze at 1 AM.

The Recycler began preparing to stash at 1:40 AM. The stash ended and Pbar resumed stacking around 2:15 AM. The Meson MTest and MCenter experiments resumed taking beam.

At 2:58 AM, Operations reported problems with a Booster quadrupole power supply (B:QL1). Experts were notified.

Operations put the TeV into shot setup at 5:15 AM and established store 4653 at 7:13 AM. The eighth transfer of antiprotons was lost. The resulting store luminosity was only 38.88E30. MTest and MCenter resumed taking beam.

The Recycler began preparing to stash at 9:26 AM. The stash ended and Pbar resumed stacking around 10:12 AM. MTest and MCenter resumed taking beam.

At 12:17 PM, MI suffered a system ground fault. This caused the beam valves to close, which resulted in the loss of the stash. Operations called in experts who found a problem at MI20. They had it repaired by 3:54 PM.

Pbar resumed stacking at 5:38 PM. NuMI and Switchyard (SY) beam also resumed.

The Recycler began preparing to stash at 8:02 PM. The stash ended and Pbar resumed stacking around 8:59 PM. MTest and MCenter resumed taking beam.

Sunday February 19

The midnight shift began with Operations monitoring store 4653 and the stash, with Pbar stacking, and with NuMI, MiniBooNE, MTest, and MCenter taking beam.

The Recycler began preparing to stash at 12:12 AM. The stash ended and Pbar resumed stacking around 12:45 AM. MTest and MCenter resumed taking beam.

The Recycler began preparing to stash at 4:25 AM. The stash ended and Pbar resumed stacking around 4:59 AM. MTest and MCenter resumed taking beam.

Operations terminated store 4653 at 6:04 AM and then put the TeV into shot setup at 6:19 AM.

Operations established store 4654 at 9:12 AM with an initial luminosity of 69.6E30.

The Recycler began preparing to stash at 10:49 AM.

MTest experimenters asked to have their beam turned off at 10:59 AM.

The stash ended and Pbar resumed stacking around 11:38 AM. MCenter resumed taking beam.

At 2:24 PM, Operations reported that MiniBooNE was off due to data acquisition (DAQ) problems. This morning, they had been off for about three hours due to the same problem.

The Recycler began preparing to stash at 3:49 PM.

MiniBooNE reported around 4 PM that they had resolved their DAQ problems.

The stash ended and Pbar resumed stacking around 4:54 PM. MCenter resumed taking beam.

MTest resumed taking beam at 7:49 PM.

The Recycler began preparing to stash at 8:57 PM. The stash ended and Pbar resumed stacking at 9:34 PM. MTest and MCenter resumed taking beam.

At 9:36 PM, the NuMI Horn tripped off for the first of many times in twenty-four minutes. Operations contacted an expert.

At 11:19 PM, Operations reported a problem with the Booster chopper. Experts were called in.

Monday February 20

The midnight shift began with Operations monitoring store 4654 and the stash, with Booster chopper problems, with Pbar stacking, with NuMI off due to Horn trouble, and with MiniBooNE, MTest, and MCenter taking beam.

At 12:10 AM, Operations turned off the Booster beam for about forty-five minutes while experts repaired the chopper.

At 12:36 AM, a Booster power supply (DOGL3) tripped off for the first of nine times over the rest of the shift. It reset each time.

Operations resumed sending beam to NuMI at 1:26 AM.

The Recycler began preparing to stash at 3:17 AM. The stash ended and Pbar resumed stacking around 3:45 AM. MTest and MCenter resumed taking beam.

MTest turned off for the day at 4:58 AM.

The Recycler began preparing to stash at 7:35 AM. The stash ended and Pbar resumed stacking around 7:57 AM. MCenter resumed taking beam.

Operations terminated store 4654 at 7:59 AM.

The Plans for Monday

The plans for today are to allow D0 and CDF to access their collision halls. After the access, TeV experts will conduct a Tevatron check and then Operations will put the TeV into shot setup.

 
Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 02/20/2006    email Fermilab